# Context Matters: Pushing the Boundaries of Open-Ended Answer Generation with Graph-Structured Knowledge Context

### 23 Jan 2024 v1

@@ 1 Introduction
In the field of artificial intelligence, LLMs Roberts et al. (2020); Kaplan et al. (2020) stand out as significant breakthroughs. They highlight the advanced capabilities of machines in both text understanding Lian et al. (2023) and generating Wei et al. (2023a). While LLMs usually produce high-quality outputs they often struggle in areas with limited resources Chen et al. (2023); Guu et al. (2020a). They also have limitations due to knowledge cutoff dates. In many cases, they hallucinate McKenna et al. (2023) when faced with very specific, low-resource or domain specific topics. There is also a trade-off between quality and the billions of parameters Xu et al. (2023), which makes the LLM challenging to explore and analyze for researchers with limited resource capacity. Recently, to address these challenges, grounding of LLMs  [^1] https://techcommunity.microsoft.com/t5/fasttrack-for-azure/grounding-llms/ba-p/3843857 and the Retrieval-Augmented Generation (RAG) method Yu et al. (2023) have been proposed. This approach allows large models to actively utilize the information from external databases as they operate. This feature is particularly useful as it means the models can provide answers that are not just based on their original training data, but are also detailed, current, and well-informed by accessing a wider range of information sources. This opens up the potential for even more factually accurate and contextually relevant responses in real-time.
@@ 2 Related Work
Large language models: Recent progress in LLMs Naveed et al. (2023); Douglas (2023) has positioned them as dominant successors to earlier transformer-based models for tackling generation tasks in natural language. Over the past several years, there has been a surge in introducing LLMs  [^7] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard like GPT3 Brown et al. (2020), Llama2 Touvron et al. (2023a) etc. for this purpose. The newly introduced LLMs have demonstrated the ability to grasp and interpret context, subsequently producing human like responses. These models generate fluent and high quality answers very well for general questions Yang et al. (2023c). So, these models are often used as the base architecture for various generation based tasks such as long Q&A Saad-Falcon et al. (2023); Kamalloo et al. (2023), multiple choice Q&A Robinson et al. (2023), dialogue systems Deng et al. (2023); Hudeček and Dušek (2023); Valvoda et al. (2022); Snell et al. (2022), chatbot systems Lee et al. (2023a); Wei et al. (2023b). However, despite the high capability of these methods, they generate factually incorrect Shuster et al. (2021) responses and are oftentimes not aligned with the actual question for domain specific knowledge. Generally, this issue of LLMs is called hallucination Dhuliawala et al. (2023); Mündler et al. (2023). Due to the cutoff knowledge, models sometime fail to provide faithful answer for a general question. To tackle this problem, recent studies Shuster et al. (2021) proposed a method of augmenting retrieved knowledge from diverse sources as a context to LLMs.
Retrieval augmented generation: Retrieval-augmented generation (RAG) Manathunga and Illangasekara (2023); Ram et al. (2023) is a framework that merges the capabilities of large pretrained language models with the advantages of external retrieval or search mechanisms. Earlier, augmenting language models with retrieved knowledge had shown effective performance for knowledge-intensive tasks Guu et al. (2020b). It is particularly useful in scenarios where a language model needs to pull in specific factual information from an extensive corpus to generate relevant and accurate outputs. In recent studies, the researchers attempt to augment external knowledge through knowledge graph prompting Wang et al. (2023), few shot domain adaptation Krishna (2023), retrieval-generation synergy Shao et al. (2023), knowledge graph-based subgraph retrieval augmented generation Kang et al. (2023), augmented adapter retriever (AAR) Yu et al. (2023).
@@ 5 Experimental setup
Evaluation metrics: We have used three metrics – ROUGE score [^15] https://huggingface.co/spaces/evaluate-metric/rouge, BERT score Zhang* et al. (2020a); Zhang et al. (2020b) and FactSumm score Heo (2021) for automatic evaluation of generated answers. Note that the FactSumm Heo (2021) package extracts the facts from the generated text and the ground truth text and computes an overall score based on the fact overlap and fact mismatch. This package has also been used in earlier works Liu et al. (2021); Qian et al. (2023) to measure the factual accuracy of the generated text.
@@ 5 Contextual Content Crux
Retrieved Content: A post briefly mentioning the XYZ error with a comment saying, ”Ï also faced this.”
@@ A.2 Pre-LLM era baselines
A.2.1 AnswerBot
Authors of this work Xu et al. [2017]; Cai et al. [2019] proposed an approach called AnswerBot, where the task is to generate a summary from diverse answers for a query. They followed three major steps – relevant question retrieval, useful answer paragraph selection, diverse answer summary generation. For retrieval, they used word2vec model and relevance calculation algorithm. For answer paragraph selection, they used various query, paragraph and user related features – relevance to query,  entity overlap,  information entropy,  semantic pattern,  format patterns,  paragraph position,  vote on answer. In answer summary generation stage, they used maximal marginal relevance (MMR) algorithm to select a subset of answer paragraphs. Further they used selected answer paragraphs to form the answer summary.
A.2.2 GenQA
Authors of this paper Hsu et al. [2021] proposed a framework to generate answers from the top candidates of a set of answer selection models. Instead of selecting the best candidates, they train a sequence to sequence transformer model to generate an answer from candidate set.
A.2.3  TechSumBot
Authors of this paper Yang et al. [2023a] show that developers frequently turn to StackOverflow for solutions, but they often encounter redundant or incomplete results. Current tools designed to summarize StackOverflow answers have clear drawbacks: they predominantly depend on manually-designed features, they struggle to filter out repetitive content, and they usually target specific programming languages. This tool autonomously produces answer summaries by extracting and ranking answers for their relevance, measuring the core importance of each sentence, and eliminating redundant details. Presented in a search engine format, TechSumBot’s efficiency is benchmarked against existing StackOverflow summary methods.
@@ A.3 Ablation study
    Based on one-shot: When focusing on one-shot methodologies, the GraphGen model achieves marginally better scores compared to the TextGen model. This suggests that the incorporation of graph structures potentially aids in better knowledge retrieval
@@ A.4 Sample prompt and generated answer
1. Archive Manager: Built-in tool. Just right-click the ISO, hit ”Open With Archive Manager”, and boom, you’re in.
@@
1. Archive Manager: Built-in. Right-click the ISO, select ”Open With Archive Manager”, and extract.

### 05 Mar 2024 v2

@@ 1 Introduction
In the field of artificial intelligence, LLMs Roberts et al. (2020); Kaplan et al. (2020) stand out as significant breakthroughs. They highlight the advanced capabilities of machines in both text understanding Lian et al. (2023) and generating Wei et al. (2023). While LLMs usually produce high-quality outputs they often struggle in areas with limited resources Chen et al. (2023); Guu et al. (2020b). They also have limitations due to knowledge cutoff dates. In many cases, they hallucinate McKenna et al. (2023) when faced with very specific, low-resource or domain specific topics. There is also a trade-off between quality and the billions of parameters Xu et al. (2023), which makes the LLM challenging to explore and analyze for researchers with limited resource capacity. Recently, to address these challenges, grounding of LLMs  [^1] https://techcommunity.microsoft.com/t5/fasttrack-for-azure/grounding-llms/ba-p/3843857 and the Retrieval-Augmented Generation (RAG) method Yu et al. (2023) have been proposed. This approach allows large models to actively utilize the information from external databases as they operate. This feature is particularly useful as it means the models can provide answers that are not just based on their original training data, but are also detailed, current, and well-informed by accessing a wider range of information sources. This opens up the potential for even more factually accurate and contextually relevant responses in real-time.
@@ 2 Related Work
Large language models: Recent progress in LLMs Naveed et al. (2023); Douglas (2023) has positioned them as dominant successors to earlier transformer-based models for tackling generation tasks in natural language. Over the past several years, there has been a surge in introducing LLMs  [^7] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard like GPT3 Brown et al. (2020), Llama2 Touvron et al. (2023a) etc. for this purpose. The newly introduced LLMs have demonstrated the ability to grasp and interpret context, subsequently producing human like responses. These models generate fluent and high quality answers very well for general questions Yang et al. (2023c). So, these models are often used as the base architecture for various generation based tasks such as long Q&A Saad-Falcon et al. (2023); Kamalloo et al. (2023), multiple choice Q&A Robinson et al. (2023), dialogue systems Deng et al. (2023); Hudeček and Dušek (2023); Valvoda et al. (2022); Snell et al. (2022), chatbot systems Lee et al. (2023a); Wei et al. (2023). However, despite the high capability of these methods, they generate factually incorrect Shuster et al. (2021) responses and are oftentimes not aligned with the actual question for domain specific knowledge. Generally, this issue of LLMs is called hallucination Dhuliawala et al. (2023); Mündler et al. (2023). Due to the cutoff knowledge, models sometime fail to provide faithful answer for a general question. To tackle this problem, recent studies Shuster et al. (2021) proposed a method of augmenting retrieved knowledge from diverse sources as a context to LLMs.
Retrieval augmented generation: Retrieval-augmented generation (RAG) Manathunga and Illangasekara (2023); Ram et al. (2023) is a framework that merges the capabilities of large pretrained language models with the advantages of external retrieval or search mechanisms. Earlier, augmenting language models with retrieved knowledge had shown effective performance for knowledge-intensive tasks Guu et al. (2020a). It is particularly useful in scenarios where a language model needs to pull in specific factual information from an extensive corpus to generate relevant and accurate outputs. In recent studies, the researchers attempt to augment external knowledge through knowledge graph prompting Wang et al. (2023), few shot domain adaptation Krishna (2023), retrieval-generation synergy Shao et al. (2023), knowledge graph-based subgraph retrieval augmented generation Kang et al. (2023), augmented adapter retriever (AAR) Yu et al. (2023).
@@ 5 Experimental setup
Evaluation metrics: We have used three metrics – ROUGE score [^15] https://huggingface.co/spaces/evaluate-metric/rouge, BERT score Zhang* et al. (2020); Zhang et al. (2020) and FactSumm score Heo (2021) for automatic evaluation of generated answers. Note that the FactSumm Heo (2021) package extracts the facts from the generated text and the ground truth text and computes an overall score based on the fact overlap and fact mismatch. This package has also been used in earlier works Liu et al. (2021); Qian et al. (2023) to measure the factual accuracy of the generated text.
@@ 5 Contextual Content Crux
Retrieved Content: A post briefly mentioning the XYZ error with a comment saying, "Ï also faced this."
@@ A.2 Pre-LLM era baselines
A.2.1 AnswerBot Xu et al. [2017], Cai et al. [2019]:
Authors of this work proposed an approach called AnswerBot, where the task is to generate a summary from diverse answers for a query. They followed three major steps – relevant question retrieval, useful answer paragraph selection, diverse answer summary generation. For retrieval, they used word2vec model and relevance calculation algorithm. For answer paragraph selection, they used various query, paragraph and user related features – relevance to query,  entity overlap,  information entropy,  semantic pattern,  format patterns,  paragraph position,  vote on answer. In answer summary generation stage, they used maximal marginal relevance (MMR) algorithm to select a subset of answer paragraphs. Further they used selected answer paragraphs to form the answer summary.
A.2.2 GenQA Hsu et al. [2021]:
Authors of this paper proposed a framework to generate answers from the top candidates of a set of answer selection models. Instead of selecting the best candidates, they train a sequence to sequence transformer model to generate an answer from candidate set.
A.2.3  TechSumBot Yang et al. [2023a]:
Authors of this paper show that developers frequently turn to StackOverflow for solutions, but they often encounter redundant or incomplete results. Current tools designed to summarize StackOverflow answers have clear drawbacks: they predominantly depend on manually-designed features, they struggle to filter out repetitive content, and they usually target specific programming languages. This tool autonomously produces answer summaries by extracting and ranking answers for their relevance, measuring the core importance of each sentence, and eliminating redundant details. Presented in a search engine format, TechSumBot’s efficiency is benchmarked against existing StackOverflow summary methods.
@@ A.3 Ablation study
    Based on one-shot: When focusing on one-shot methodologies, the GraphGen model achieves marginally better scores compared to the TextGen model. This suggests that the incorporation of graph structures potentially aids in better knowledge retrieval.
@@ A.4 Sample prompt and generated answer
1. Archive Manager: Built-in tool. Just right-click the ISO, hit "Open With Archive Manager", and boom, you’re in.
@@
1. Archive Manager: Built-in. Right-click the ISO, select "Open With Archive Manager", and extract.